### v1

# When Machines Get It Wrong: Large Language Models Perpetuate Autism Myths More Than Humans Do

Abstract
As Large Language Models become ubiquitous sources of health information, understanding their capacity to accurately represent stigmatized conditions is crucial for responsible deployment. This study examines whether leading AI systems perpetuate or challenge misconceptions about Autism Spectrum Disorder, a condition particularly vulnerable to harmful myths. We administered a 30-item instrument measuring autism knowledge to 178 Spanish participants and three state-of-the-art LLMs including GPT-4, Claude, and Gemini. Contrary to expectations that AI systems would leverage their vast training data to outperform humans, we found the opposite pattern: human participants endorsed significantly fewer myths than LLMs (36.2% vs. 44.8% error rate; z=2.59z=-2.59, $p=.0048$). In 18 of the 30 evaluated items, humans significantly outperformed AI systems. These findings reveal a critical blind spot in current AI systems and have important implications for human-AI interaction design, the epistemology of machine knowledge, and the need to center neurodivergent perspectives in AI development.
Keywords: Large Language Models; Autism Spectrum Disorder; Health misinformation; Human-AI comparison; Digital epistemology; Neurodiversity
1 Introduction
The emergence of Large Language Models as primary information sources represents a fundamental shift in how the public accesses health-related knowledge [1]. With over 100 million weekly active users of ChatGPT alone, and growing integration of AI assistants into healthcare and educational platforms, the accuracy of LLM-generated health information has become a matter of significant public concern [2]. Early evidence from healthcare professionals interacting with ChatGPT reveals nuanced patterns of trust and satisfaction that vary with professional experience and medical specialty [3], while public willingness to adopt AI-based medical consultation systems remains contingent on perceived usefulness and prior satisfaction with human doctors [4]. This concern is particularly acute for conditions surrounded by persistent myths and social stigma, where misinformation can have profound consequences for affected individuals and their families.
Autism Spectrum Disorder exemplifies such a condition. Despite substantial advances in scientific understanding, autism remains shrouded in misconceptions that perpetuate stigma, delay diagnosis, and impede access to appropriate support [5]. Common myths persist in public discourse, including beliefs that autistic individuals lack empathy, cannot form relationships, or that autism results from inadequate parenting [6, 7]. These misconceptions translate into real-world discrimination, underdiagnosis particularly among women and minorities, and inadequate service provision [8]. The emergence of LLMs as epistemic authorities introduces a novel dimension to this challenge: if these systems perpetuate autism myths, they may amplify existing misinformation at unprecedented scale, whereas if they demonstrate superior accuracy to lay populations, they could serve as valuable tools for public education.
This study provides the first systematic comparison of autism knowledge between leading LLMs and a human population sample. We designed a 30-item instrument capturing both myths and facts about autism, administered it to 178 Spanish participants, and posed identical items to GPT-4, Claude, and Gemini. Our investigation examined whether LLMs demonstrate greater or lesser accuracy than humans in distinguishing autism myths from facts, whether systematic differences exist in human-AI performance gaps across categories of autism knowledge, and what human characteristics predict better autism knowledge. The findings reveal a counterintuitive pattern: despite their vast training corpora, state-of-the-art LLMs endorse autism myths at significantly higher rates than human respondents.
2 Theoretical Background
Public understanding of autism has evolved dramatically since Leo Kanner’s initial descriptions in 1943, yet each era’s conceptualizations have left residue in contemporary beliefs. The refrigerator mother theory, though thoroughly debunked, still echoes in assumptions about parental causation, while the savant stereotype, amplified by popular media, persists despite describing only a small minority of autistic individuals [6]. Contemporary autism science emphasizes heterogeneity, recognizing that autism manifests differently across individuals, contexts, and developmental stages [9]. The DSM-5 reconceptualization as a spectrum condition formally acknowledged this diversity [10], yet media representations continue to privilege narrow archetypes, rendering invisible the majority of the autistic population [7].
The empathy myth warrants particular attention. The notion that autistic people lack empathy derives from early theory of mind research suggesting difficulties in inferring others’ mental states [11]. However, subsequent research has complicated this picture considerably, with Bottema-Beutel et al. [5] demonstrating that autistic individuals experience and express empathy differently rather than deficiently. The double empathy problem proposed by Milton [12] reframes communication difficulties as bidirectional, noting that neurotypical people also struggle to understand autistic perspectives. These nuances rarely penetrate public consciousness, where the simpler deficit narrative persists.
Understanding why LLMs might perpetuate myths requires examining how they acquire and represent knowledge. Unlike humans, who learn through embodied experience, social interaction, and explicit instruction, LLMs learn statistical regularities in text [13]. Their knowledge consists of patterns reflecting which words follow which, how concepts cluster, and what phrasings recur across their training corpus. This learning mechanism carries important implications: LLMs inherit the biases of their training data, and if internet text overrepresents deficit framings of autism, LLMs will learn to associate autism with deficit [14]. Furthermore, LLMs lack the experiential grounding that enables humans to recognize when received wisdom is wrong. A person who befriends an autistic individual encounters direct evidence against the empathy myth, whereas an LLM encounters only text about empathy and autism, much of which perpetuates misconception. Recent work on LLM health knowledge has produced mixed findings, with Singhal et al. [15] reporting impressive performance on medical licensing examinations while Abd-Alrazaq et al. [16] identified significant limitations in mental health contexts where nuance and contextual sensitivity matter more than factual recall. Autism, sitting at the intersection of medical and social domains, may represent a particularly challenging case.
Spain provides a valuable context for this examination. Spanish society has invested significantly in autism awareness through government-published media guidelines [17] and active advocacy organizations, and autism prevalence and identification have increased markedly over the past decade, raising public familiarity with the condition [18]. Yet research documents persistent misconceptions even among education professionals [19], suggesting that awareness initiatives have achieved incomplete penetration.
3 Method
We employed a cross-sectional, comparative design contrasting autism knowledge between human participants and LLMs. The study received ethical approval from Universidad Pontificia Comillas Research Ethics Committee, and all human participants provided informed consent.
The autism knowledge instrument was developed through a systematic process designed to ensure content validity. Initial items were generated through literature review, extracting autism-related claims from peer-reviewed research on misconceptions [20, 5, 6], DSM-5 diagnostic criteria [10], and Spanish institutional documents [17, 19]. The final instrument contained 30 items comprising 16 myths and 14 facts. Respondents indicated agreement on a 5-point scale ranging from strongly disagree to strongly agree, and an attention check question was included to filter unreliable responses.
Human participants were recruited through social media, university networks, and community contacts during May 2025, with eligibility requiring age of at least 16 years, current Spanish residency, and Spanish language fluency. Of 200 initial responses, 178 passed attention checks and were retained for analysis. The sample was 61.8% female and 37.6% male, with 0.6% identifying as other gender. The mean age was 36.7 years (SD = 16.7) with a range from 16 to 89 years. Educational attainment was high: 60.1% held a university degree and 28.7% had postgraduate education, while 6.2% had completed high school, 2.3% vocational training, and 2.2% had secondary or primary education. Notably, 19.7% of participants reported having a close relationship with someone with autism, whether a family member, friend, or colleague. While not representative of the general Spanish population, this educated sample provides a conservative test of our hypotheses: if LLMs underperform relative to an educated sample, concerns about deployment for general populations are amplified.
Three leading LLMs were selected based on market prominence: GPT-4 from OpenAI, Claude from Anthropic, and Gemini from Google. API access was obtained through commercial subscriptions. Python scripts were developed using the official APIs to automate requests. Each of the 30 statements was sent to each model, and responses were recorded and transformed into proportions of agreement expressed as percentages. Human responses were dichotomized with ratings of 4 or 5 (agree or strongly agree) coded as believing the myth. Group differences were tested using two-proportion z-tests from the statsmodels library in Python.
4 Results
Across all 30 items, human participants demonstrated significantly lower myth endorsement rates than the pooled LLM average. The mean proportion of incorrect responses was 36.2% for humans compared to 44.8% for LLMs. This 8.6 percentage point difference reached statistical significance (z=2.59z=-2.59, $p=.0048$). Table 1 presents the comparison across respondent types.
Table 1: Overall Error Rates by Respondent Type
Respondent Error Rate %
Human participants 36.2
GPT-4 41.6
Claude 44.1
Gemini 48.7
LLM Average 44.8
Among AI systems, GPT-4 performed best with 41.6% error rate, showing knowledge relatively aligned with scientific consensus though still exceeding the human average. Claude showed similar behavior with 44.1% error rate, though with a tendency to respond more conservatively and avoid blunt statements. Gemini exhibited the highest error rate at 48.7%, suggesting it is the least reliable model on issues related to neurodivergence, possibly due to training bias towards social media content.
In 18 of the 30 evaluated items, humans showed significantly better performance than LLMs ($p<.05$). In the remaining 12 items, the difference was not significant or the AIs were more aligned with expert knowledge. The myths where humans clearly outperformed AIs included statements such as “autism completely blocks learning,” “autistic people have no empathy,” and “they cannot keep up with school.” These are myths concerning the social, emotional, and educational capacities of autistic individuals, areas where human experience and social awareness appear to provide epistemic advantages.
Conversely, the myths where LLMs offered better answers were those more technical or widely debunked in scientific literature, such as “autism is a mental illness,” “the use of screens causes autism,” and “all autistic people are the same.” This distribution suggests that LLMs are trained to detect obvious fallacies well-documented in their training data but do not always handle nuances or the diversity within the autism spectrum adequately.
The finding that 19.7% of human participants had a close relationship with someone with autism may have positively influenced human performance, as direct contact with autistic individuals provides experiential evidence against common myths that text-based learning cannot replicate.
5 Discussion
This study provides the first systematic comparison of autism knowledge between leading LLMs and a human population sample. The central finding is that humans substantially outperformed AI systems in identifying autism myths, with the overall difference reaching statistical significance. This result is counterintuitive given that LLMs are trained on vastly more text than any human could process, yet humans demonstrated superior judgment on the majority of items.
The explanation lies in the nature of knowledge itself. There are things one can learn from reading and things one can only learn from living. The myth that autistic people lack empathy is refuted in clinical literature, but it is refuted more powerfully by direct experience with autistic individuals. LLMs, by design, have access only to text; they can process millions of sentences about autism but cannot befriend an autistic person or witness their emotional expressions firsthand. The fact that nearly one in five participants reported personal connections to autistic individuals may explain part of the human advantage, as such connections provide direct evidence against prevalent stereotypes.
The pattern of results also reveals important domain differences. LLMs excelled at rejecting myths that are clearly debunked in scientific and medical literature, such as claims about screen time causing autism or autism being a mental illness. These facts are well-represented in training data with clear scientific consensus. However, LLMs struggled with myths concerning the social and emotional lives of autistic individuals, areas where knowledge is more experiential, where autistic voices have historically been marginalized in the literature, and where deficit-based framings have dominated clinical discourse.
This pattern has practical implications. Users consulting LLMs about biomedical aspects of autism may receive accurate information. Users asking about whether autistic individuals can feel love, want friendships, or experience empathy may receive responses that perpetuate harmful misconceptions. For designers of AI-powered health resources, these findings suggest the need for domain-specific validation, transparent disclosure of limitations, and mechanisms to direct sensitive queries to human experts or community resources. Notably, Isaac et al. [21] showed that making human cognitive biases salient can increase acceptance of medical AI, suggesting that framing AI as a complement rather than a replacement may be effective—but our results caution that such framing must be accompanied by rigorous validation of AI accuracy in socially sensitive health domains.
The differences among LLMs also merit attention. Gemini’s substantially higher error rate compared to GPT-4 suggests that model architecture, training data curation, and safety tuning all influence autism knowledge. These differences could inform deployment decisions and highlight the importance of model-specific evaluation rather than treating all LLMs as equivalent.
Several limitations should be acknowledged. Our human sample was notably well-educated, which limits generalizability but provides a conservative test of our hypotheses. LLM capabilities evolve rapidly, and our findings represent a snapshot of May 2025 model versions. Future research should examine how LLM performance changes over time and whether targeted interventions in training can address identified weaknesses. Most importantly, research should center autistic perspectives, involving autistic individuals not merely as subjects but as experts and co-investigators.
6 Conclusion
This study demonstrates that state-of-the-art Large Language Models perpetuate autism myths at rates significantly exceeding those of human populations, with humans outperforming LLMs in 18 of 30 evaluated items. While LLMs excel at rejecting well-documented biomedical misconceptions, they struggle with myths concerning the social and emotional dimensions of autism. These findings reveal fundamental limitations in how current AI systems represent human neurodiversity and carry urgent implications for AI governance and deployment. For the millions of users who consult AI about autism, current systems may reinforce rather than correct the stigma that harms autistic individuals. The humans in our sample, particularly those with personal connections to autistic individuals, demonstrated knowledge that billions of parameters could not replicate: that autistic people are fully human, fully feeling, and fully worthy of understanding and connection.
Declaration of Interest
None.
CRediT Author Statement
Eduardo C. Garrido-Merchán: Conceptualization, Methodology, Software, Formal analysis, Writing – Original Draft, Supervision. Adriana Constanza Cirera Tirschtigel: Investigation, Data curation, Writing – Review & Editing.
Data Availability
Data and analysis code are available upon request.